Calculate the specific heat of a substance if 420 calories is needed to raise the temperature of

this 75 grams substance from 15 ᵒC to 50 ᵒC.



1
Expert's answer
2021-05-25T10:10:03-0400

The heat required to change the temperature:


"Q=cm\Delta t,\\\space\\
c=\frac{Q}{m\Delta t}=\frac{420}{75·(50-15)}=0.16\text{ cal/(g°C)}."